40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 170 of 349  Not logged in ELOG logo
ID Date Author Type Category Subject
  9044   Wed Aug 21 00:18:03 2013 MasayukiSummaryGreen LockingX-arm PDH OLTF measurement

[Manasa Masayuki]
Today we measured the openloop transfer function of the PDH green lock of the x-arm.

Edit //manasa// The excitation was given from SR785 source. SR560 was used as the summing node at the PDH servo box output where the loop was broken to measure the OLTF. The SR785 was used to measure the frequency response (CH2/CH1; CH1 A SR560 output and CH2 A PDH servo output) in sweptsine mode.

We measured with two different servo gain. We started with the servo gain of 3 and at that gain the UGF was 1.5 kHz and the phase margin was 50 degree. After that we increase the servo gain to 5.5 and at that gain the UGF was 6.2 kHz and the phase margin was 55 degree. In all the measurement we use the source amplitude of 1.0 mV for all frequencies (from 100 Hz to 100 kHz). We could not increase the gain and also the source amplitude any more because the green was kicked out of lock.

Next work list
1. In the earlier measurements we found the UGF of the PDH green lock of the x-arm as 10 kHz and the phase margin as 45 degree, so we will investigate what has changed from these measurements.elog 4490

2. We will measure the power spectrum of the error signal and the feedback signal.

3. We will calibrate the above signals to compare with ALS out of loop noise.

netgpib was taking forever to transfer data. So the measurements are just photos of the display.

attachment1 - servo gain 3

IMG_1226.JPG

attachment2 - servo gain 5.5

IMG_1228.JPG

  9043   Tue Aug 20 18:42:57 2013 JenneUpdateLSCREFL investigations

I have done the swap in the REFL path.  First, I swapped the positions of REFL11 and REFL55.  Then, I swapped out the 50/50 BS for a 90% reflection BS.  (90% goes to REFL55, 10% goes to REFL11).  I also changed the aluminum dump that was dumping the old REFL165 path into a razor dump.

Before: REFL11 had 4.0mW, REFL55 had 3.1mW.  Now, REFL11 has 0.53mW, and REFL55 has 6.9mW.  REFL165 still has around 61mW of light, and REFL33 has 3.3mW (the things that were changed were after 165 and 33 in the REFL path). 

Now, the DC value of the REFL PDs are:  REFL165 = 10.4V, REFL33 = 110mV, REFL55 = 232mV, REFL11 = 18.6mV. 

As I was finishing aligning the beams onto all of the REFL diodes, Manasa asked for the IFO so she and Masayuki could continue their work on the Xarm, so I'll check the signals acquired a little later.

  9042   Tue Aug 20 16:23:41 2013 ranaSummaryGeneral/home/cds nearly full

/home/cds is >98% full - below are some of the usage numbers:

controls@rosalba:/users/OLD 0$ du -h --max-depth=1
42M    ./katrin
1.5M    ./ben
2.4M    ./sanjit
569M    ./waldman
328M    ./sonia
3.6G    ./lsinger
44M    ./dbusby
105M    ./dbarron
21M    ./manuel
709M    ./yaakov
46M    ./rodionov
240M    ./ishwita
2.7G    ./clara
56M    ./gopal
290M    ./mashaB
87M    ./varvella
5.6M    ./Sascha
2.9G    ./ryan
190M    ./nancy
3.5G    ./john
269M    ./elizabeth.davison
165M    ./jweiner
460K    ./mjones
49M    ./stephanie
52M    ./mohana
56M    ./noriyasu
38M    ./mjenson
76M    ./sballmer
224M    ./kirk
812K    ./bonnie
33M    ./janosch
16M    ./kevin
122M    ./dblair
2.6G    ./mirko
389M    ./keenan
195M    ./tf
150M    ./littlezach
193M    ./jmiller
1.8G    ./ting
131M    ./dmalling
842M    ./sharmila
1.4G    ./caryn
12G    ./rward
4.1M    ./jay
443M    ./emintun
184M    ./katharine
76K    ./nick
804K    ./nicole.ing
14M    ./jenny
542M    ./vsanni
45M    ./peter
7.8G    ./miyakawa
4.8M    ./channa
4.0K    ./frank
9.9G    ./razib
35M    ./amin
361M    ./sharon
62M    ./bram
3.9M    ./volodya
7.9M    ./larisa
301M    ./sasha
33M    ./eric.hendries
18M    ./vuk
101M    ./huan
1.8M    ./sonali
453M    ./megan
43M    ./Royal
5.4G    ./ayaka
19M    ./mott
518M    ./justing
501M    ./avi
173M    ./kakeru
3.9G    ./alberto
41M    ./paul.fulda
59M    ./elena
67G    .

controls@rosalba:/opt/rtcds/userapps 0$ du -h --max-depth=1
1.4G    ./tags
13M    ./trunk.bak
40K    ./.svn
3.0G    ./trunk
174M    ./trunk.bak2
4.2G    ./branches
8.7G    .

linux1:cds>nice du -h --max-depth=1
du: `./llo/chans/daq/archive': Permission denied
du: `./llo/chans/daq/old': Permission denied
707M    ./llo
9.7M    ./mit~
752K    ./raidwebFirmware
462M    ./epics
2.1G    ./tmp
1.5G    ./gds
76M    ./project
9.1G    ./ligo
449G    ./rtcds
3.3G    ./apps
20K    ./.kde
512K    ./cdscfg
1.4M    ./.Trash-controls
5.8M    ./scripts
20K    ./.TemporaryItems
964G    ./caltech
71M    ./bin
16K    ./.Trash-1001
4.5G    ./rtapps
564M    ./src
11M    ./vw
3.8M    ./dvSave
460M    ./lho
1.2G    ./data
1.5T    .

  9041   Tue Aug 20 11:52:20 2013 JenneUpdateLSCREFL investigations

Quote:

As I always tell everyone: Don't use a 10% reflector which produce ghost beams. Use a 90% reflector. 

 Hmmm, yes, I forgot (bad me).  I'll find a 90% refl BS, and swap the positions of REFL11 and REFL55.

  9040   Tue Aug 20 11:41:30 2013 KojiUpdateLSCREFL investigations

As I always tell everyone: Don't use a 10% reflector which produce ghost beams. Use a 90% reflector.

  9039   Tue Aug 20 10:59:15 2013 SteveUpdateGreen LockingXend green layout corrections

Quote:

 Shutter moved, no more clipping.

Pick-off mirror 2" replaced by 1" one. Laseroptik HR 532nm, incident angle 30-45 degrees, AR 532 nm

Green REFL PD moved to 4" close to pick-off mirror. Pd being close to pick-off does not separate multiple reflections on it. I'll replace Laseroptic mirror with Al one. It is not easy to find.

 Hole cut into side wall for doubler oven cable to exit.

 

 

 Beam trap for Pd refl is in place. Cabeling is ti·died up.

 Laseroptic 1" mirror is replaced by Al 1" mirror. Problem remains the same. This diffraction patter has to be coming from the Faraday.

  Atm1, good separation when Pd is far 

  Atm2, bad separation when Pd is close 

Attachment 1: faraway.jpg
faraway.jpg
Attachment 2: closer.jpg
closer.jpg
  9038   Tue Aug 20 01:28:47 2013 JenneUpdateLSCREFL investigations

According to the wiki, REFL 11 has a transimpedance of 4.08kV/A, and REFL 55 has a transimpedance of 615V/A.  This is a ratio of ~6.5 .  My optickle simulations from earlier this evening indicate that, at maximum, there is a ~factor of 2 more signal in REFL 11 than REFL 55.  This is a factor of order 10-15.  Then, REFL 55 has 15dB whitening gain, which is a factor of ~4.  So, this explains why we're seeing so much more digital signal on REFL11 than REFL55.

Tomorrow, I need to replace the 50/50 beam splitter that splits the beam between REFL55 and REFL11 (33 and 165 have already had their light picked off at this point).  I want to put in a 10% reflector, 90% transmission beamsplitter.  Steve, can you please find me one of these, and if we don't have one, order one? This will give us a little more light on 55, and less light on 11, so hopefully we won't be saturating things anymore.

 

  9037   Tue Aug 20 00:19:23 2013 ranaUpdateLSCPRMI / DRMI investigations

While Jenne was plotting, I locked and aligned the MICH with AS55_Q. Then I aligned the PRM and locked PRMI using REFL55_I/Q with triggering on POP22, but no power normalization.

I used this to set the phase for REFL11 and REFL55 (driving PRM at 111.3 Hz and minimizing the Q response using the DTT Sine Response tool). I flipped the sign on REFL11 by 

The REFL11 gain is ~50x larger than REFL55; this is with the 15 dB whitening gain on REFL55 and none for REFL11. What's going on here? The attached PDF shows the two time series with the free swinging PRMI and both phases set to ~ +/- 2 deg. The REFL55 signals have been scaled up by 50x.

So then we went in and looked at the RF signals at the demod boards. To do this we disconnected the RFPD test cables and hooked the RF Mon outputs into the 50 Ohm inputs on a scope. The following PNG images show the scope traces. The REFL11 (yellow) traces are too big!! See how small the REFL55 (green) are. REFL11 is saturating - need to fix.

TEK00000.PNGTEK00001.PNG

TEK00002.PNGTEK00003.PNG

Attachment 1: REFL.pdf
REFL.pdf
Attachment 6: REFL-2.pdf
REFL-2.pdf
  9036   Mon Aug 19 23:08:31 2013 JenneUpdateLSCDRMI sensing signals

Here are a bunch of sensing signals.  The configuration is always DRMI.  Except for the optic noted in the title and the x-axis of any individual plot, other optics are held in their nominal position.  DRMI condition is sidebands resonant in PRCL, 55MHz sideband resonant in SRCL.  Each plot has an error signal, as well as the 2f signals at POP and AS.

 The phases of POP22 and POP110 have been adjusted so that the I signal is maximized when everything is at the nominal positions (sideband resonant for PRMI).  The phase of AS110 has been adjusted so that the I signal is maximized when the DRMI is in the nominal position (f2 resonant in SRC).  The phases of the 1f1, 1f2, 2f1 and 2f2 REFL signals were all adjusted to have max PRCL signal in the I phase.  AS55 was adjusted to have max SRCL signal in the Q phase.

 DRMI_PRM_REFL11I.png

DRMI_PRM_REFL33I.png

DRMI_PRM_REFL55I.png

DRMI_PRM_REFL165I.png

DRMI_MICH_REFL11Q.png

DRMI_MICH_REFL33Q.png

DRMI_MICH_REFL55Q.png

DRMI_MICH_REFL165Q.png

DRMI_SRM_REFL55Q.png

DRMI_SRM_REFL165Q.png

DRMI_SRM_AS55Q.png

 

  9035   Mon Aug 19 19:08:35 2013 KojiUpdateGreen LockingXend green layout corrections

- An Aluminum mirror instead of 2" unknown mirror for the pick-off for the rejected beam from the green faraday isolator (Steve)
=> Replaced. To be reviewed

- Faraday mount replacement. Check what we have for the replacement. (Steve)

- The green REFL PD should be closer to the pick-off mirror. (Steve)
=> Moved. To be reviewed

- A beam dump should be placed for the green REFL PD

- Move the green shutter to the place where the spot is small (Steve)
=> Moved. To be reviewed.

- The pole of the PZT mounting should be replaced with a reasonable one. (Steve with Manasa's supervision)

- Tidying up doubling oven cable. Make a hole on the wall. (Steve)
=> Done. To be reviewed.

- Tidying up the PZT cabling (Steve)

- The optics are dirty. To be drag wiped. (Manasa, Masayuki)

  9034   Mon Aug 19 17:40:32 2013 SteveUpdateGreen LockingXend green layout corrections

 Shutter moved, no more clipping.

Pick-off mirror 2" replaced by 1" one. Laseroptik HR 532nm, incident angle 30-45 degrees, AR 532 nm

Green REFL PD moved to 4" close to pick-off mirror. Pd being close to pick-off does not separate multiple reflections on it. I'll replace Laseroptic mirror with Al one. It is not easy to find.

 Hole cut into side wall for doubler oven cable to exit.

 

 

Attachment 1: beforeC.jpg
beforeC.jpg
Attachment 2: nowC.jpg
nowC.jpg
Attachment 3: stillMultiple.jpg
stillMultiple.jpg
  9033   Mon Aug 19 16:18:56 2013 manasaUpdateGreen LockingXend green aligned

ASX scripts for PZT dither have been fixed appropriately. Script resides in scripts/ASX.

You can run the scripts from the ASX medm screen now.

  9032   Mon Aug 19 15:23:07 2013 KojiUpdateIOOMC mirrors' ASC has non-zero inputs

[Jenne, Koji]

This disturbance in the MC ASC channels were fixed.

This craziness happened ~10pm last night. Was there any action at the time? >> Sunday-night workers? (RXA: No, Nakano-kun and I left before 9:30 PM)

We found that the signals came from c1ioo. However, restarting, recompiling c1ioo and c1mcs didn't help
to clean up this issue. Just in case we cleaned up the corresponding entries in the ipc file /opt/rtcds/caltech/c1/chans/ipc/C1.ipc
and recomplied c1ioo and c1mcs because these are the channels we touched last week to mitigate the timing out issue of c1rfm.

Incidentally, we fell into a strange mode of the RCG: IOPs could not restart. We ended up running "sudo shutdown -r now"
on each machine (except for c1lsc which was not affected by this issue). This solved the issue.

Even now c1oaf could not be running properly. This is not affecting the IFO operation right now, but we need to look into this issue again
in order to utilize OAF.

  9031   Mon Aug 19 14:22:36 2013 ranaUpdateGreen LockingXend green aligned

  9030   Mon Aug 19 11:30:20 2013 JenneUpdateIOOMC mirrors' ASC has non-zero inputs

[Masayuki, Jenne]

When I came in this morning, I noticed that the Mode Cleaner had not been locked for at least the past 8 hours.  We moved the MC SUS sliders until the MC SUSPIT and SUSYAW values for each mirror were back to approximately the place they were the last time the MC was nicely locked (~12 hours ago).  This got the MC flashing TEM00, so we thought we were doing well. 

However, if the servo was enabled, any time the cavity flashed a small-order mode (especially 00), the mirrors would get super kicked.  Not good

We went to investigate, and discovered that the RFPD aux laser was left on again.  We turned that off, however that didn't fix the situation. 

Manasa suggested checking that the WFS were really, really off.  When we looked at the WFS master screen, we noticed that although the WFS servos were off, the MC mirrors' ASC filter banks had non-zero inputs.  We checked, and this is not from the MCASS, nor is it from the MC WFS lockins.  At this point, I have no idea where these signals are coming from.  I have turned off the ASC outputs for all the MC mirrors (which means that we cannot turn on the WFS), and the MC locks fine

So, we need to know where the ASC signals are coming from.  There isn't anything that I can see, from any screen that I can find, that indicates some signals being sent over there.  Has anyone done anything lately?  I know Koji was working on IPC stuff the other day, but the MC was locking fine over the weekend until yesterday afternoon, so I suspect that's not the culprit. 

I have turned off the outputs of the WFS lockins, as part of my turning things off, so if whatever script needs them doesn't enable them, they should be turned back on by hand.

  9029   Mon Aug 19 11:12:54 2013 SteveUpdateVACRGA scan at day 13

 

 

Attachment 1: rgaScan13d.png
rgaScan13d.png
  9028   Mon Aug 19 10:16:15 2013 PicassoMetaphysicsTreasureoutsider art

 ranasglory.png

  9027   Mon Aug 19 10:03:17 2013 SteveHowToGeneralhow not to leave a cable

We can not leave cables connected like this. This is a burned toast award.

Attachment 1: NONOcable.jpg
NONOcable.jpg
  9026   Mon Aug 19 09:54:13 2013 SteveUpdatesafetyMasayuki receives safety training

Masayuki Nakano, a student of Seiji's from ICRR / U Tokyo, is visiting us here at the 40m lab for the next couple months.

He received 40m specific basic safety training this morning.

  9025   Mon Aug 19 09:36:32 2013 KojiUpdateGreen LockingXend green aligned

[Rana Koji]

This is an elog about the activity on Friday night.

- The X arm green beam was aligned with assist of the ASX system.

- M1 PZT alignment was swept while M2 PZT was under the control of ASX.

- Everytime M1 was touched, M2 was restored by manual alignment so that the REFL beam hits the center of the REFL PD.
  This way we could recover the lock of TEM00. Once TEM00 is recovered, ASX took care of the alignment of M2

- The error signal used by the cavity dither did not give us a good indication where the optimal alignment is.

- Thus the best alignment of M1 had to be manually scanned. The resulting maximum green transmission was ~0.88

- Once the beam was aligned, the out-of-loop stability of the Xarm was measured.
  There has been no indication of the improvement compared to Manasa's measurement taken before our beam alignment.

Attachment 1: ALS_OUTOFLOOP_130816.pdf
ALS_OUTOFLOOP_130816.pdf
  9024   Mon Aug 19 07:53:48 2013 SteveUpdateSUSETMX damping restored

ETMX sus damping restored

  9023   Sun Aug 18 20:07:41 2013 ranaUpdateComputer Scripts / Programsuserapps SVN up

JoeB and JamieR are working somewhat coherently on a set of python libraries to fulfill all of our command line CDS wants. This is being done mostly to satisfy The Guardian and the SkunkTools project.

I did an 'svn up' in /opt/rtcds/userapps (it might finish in ~1000 years) to get the things that they have so far (in particular, Joe's 'pyavg'). There's going to be some issues since the pylib stuff written by Yuta/Kiwamu has never been integrated with anything and is imported as 'epics' in many python scripts. As we move over to the new stuff there will be a lot of broken script functions since the new libraries are also used in that way.

  9022   Sun Aug 18 17:56:16 2013 ranaSummaryCDSMEDM Screen CPU Usages

I noticed at LLO (?) that the LSC screen there uses up ~25-30% of the CPU time on a single core for the control room iMac workstations - this seems excessive.

Here is an accounting of CPU usage percentages for some of our screens:

 

Screen Name CPU (%)
LSC_OVERVIEW 7
ALS_OVERVIEW 0
ALS 1
SUS_SUMMARY 0
IOO_WFS_MASTER 0.3
OPLEV_MASTER 0.5

These were measured using the program 'glances' on rosalba. MEDM running with only the sitemap used up 0.9% of a CPU. With the screens running, the fluctuation from sample to sample could be ~ +/- 0.5%. While the LSC screen seems to be the biggest pig, it is only big in comparison to small pigs. Certainly this pig has gotten bigger after getting sent to Louisiana.

Attachment 1: obama1404_666531c.jpg
obama1404_666531c.jpg
  9021   Sun Aug 18 16:04:07 2013 ranaSummaryCDSFB lights all RED: mxstream restart

Sun Aug 18 15:52:50 2013

Found the FB lights (C1:FEC-NN_FB_NET_STATUS and C1:DAQ-DC0_C1XXX_STATUS) RED for everything on the CDS_FE_STATUS screen.

I used the (! mxstream restart) button ro restart the mxstreams. Everything is green now.

PMC was out of lock- relocked it and the IMC locked itself as did the X & Y arms on IR. X was already green locked.

Attachment 1: IFO-Trend.png
IFO-Trend.png
  9020   Fri Aug 16 21:15:04 2013 ranaUpdateCDSNew/old CDS laptop for X-End

I took the "aso-laptop" and made it into Ubuntu a couple months ago. Today I added it to the Martian network and then moved it to the X End.

I followed the instructions in (https://wiki-40m.ligo.caltech.edu/Network) and added it to the files in /var/named/chroot/var/named on linux1 and did the "service named restart".

The router already had his MAC address in its list (because Yoichi was illegally using his personal laptop on the Martian). The new laptop's name is 'asia'. This is a legal name according to our computer naming conventions and this Wikipedia page (http://en.wiktionary.org/wiki/Category:Italian_female_given_names). It has been added to the Name Pool on the wiki.

The terminal on the laptop still calls itself 'aso-laptop' so I need some help in fixing that. It successfully connects to 40MARS and displays a MEDM sitemap after sshing in to pianosa.

I use 'ssh -X -C' since I find that compression actually helps when the laptops are so far from the router.

  9019   Fri Aug 16 19:36:49 2013 CharlesUpdatePSLPMC_trans Channel

Rana and I connected the PMC_trans output to the BNC connector board on the west end of the PSL table (the channel is labeled). I took a few spectra off of PMC_trans and the SR785 was connected directly to the PMC_trans output for about an hour.

Data will follow.

  9018   Fri Aug 16 13:25:50 2013 KojiUpdateASSASX model/screen cleaning up

[Koji Manasa]

Yesterday we cleaned up the ASX model and screens to have more straight forward structure of the screen
and the channel names, and to correct mistakes in the model/screens.

The true motivation is that I suspect the excess LF noise of the X arm ALS can be caused by misalignment
and beam jitter coupling to the intensity noise of the beat. I wanted to see how the noise is affected by the alignment.
Currently X-end green is highly misaligned in pitch.

- Any string "XEND" was replaced by "XARM", as many components in the system is not localized at the end table.

- The name like "XARM-ITMX" was changed to "XARM-ITM". This makes easier to create the corresponding model for the other arm.

- There was some inconsistency between the MEDM screens and the ASX model. This was fixed.

- A template StripTool screen was created. It is currently saved in users/koji/template as ASX.stp.
  It will be moved to the script directory once it's usefulness is confirmed.


The next step is to go to the end table and manually adjust M2 mirror while M1 is controlled by the ASX.
The test mass dithering provides the error signal for this adjustment but the range of the PZT is not enough
to make the input spot position to be controlled. In the end, we need different kind of matching optics
in order to control the spot position. (But is that what we want? That makes any PZT drift significantly moves the beam.)

  9017   Fri Aug 16 09:35:18 2013 SteveUpdateVACVacuum Normal state recognition is back

Quote:

Quote:

Quote:

Quote:

Apparently all of the ION pump valves (VIPEE, VIPEV, VIPSV, VIPSE) opened, which vented the main volume up to 62 mTorr.  All of the annulus valves (VAVSE, VAVSV, VAVBS, VAVEV, VAVEE) also appeared to be open.  One of the roughing pumps was also turned on.  Other stuff we didn't notice?  Bad. 

 Several of the suspensions were kicked pretty hard (600+ mV on some sensors) as a result of this quick vent wind.  All of the suspensions are damped now, so it doesn't look like we suffered any damage to suspensions.

CLOSE CALL on the vacuum system:

Jamie and I disabled V1, VM2 and VM3 gate valves by disconnecting their 120V solenoid actuator before the swap of the VME crate.

The vacuum controller unexpectedly lost control over the swap as Jamie described it. We were lucky not to do any damage! The ion pumps were cold and clean. We have not used them for years so their outgassing possibly  accumulated to reach ~10-50 Torr

I disconnected_ immobilized and labelled the following 6 valves:  the 4 large ion pump gate valves and VC1,  VC2  of the cryo pump. Note: the valves on the cryo pump stayed closed. It is crucial that a warm cry pump is kept closed!

This will not allow the same thing to happen again and protect the IFO from warm cryo contamination.

The down side of this that the computer can not identify vacuum states any longer.

This vacuum system badly needs an upgrade. I will make a list.

 While I was doing the oil change of the roughing pumps I accidentally touched the 24 V adjustment knob on the power supply.

All valve closed to default condition. I realized that the current indicator was red at 0.2A  and the voltage fluctuated from 3-13V

Increased current limiter to 0.4A and set voltage to 24V     I think this was the reason for the caos of valve switching during the VME swap.

 

 Based on the facts above I reconnected VC1 and VC2 valves.  State recognition is working.  Ion pumps are turned off and their gate valves are disabled. 

We learned that even with closed off gate valves while at atmosphere  ion pumps outgass hydrocarbons at 1e-6 Torr level.  We have not used them for this reason in the passed 9 rears.

 

I need help with implementing V1 interlock triggered by Maglev failure signal  and-or P2 pressure.

MEDM screen agrees with vacuum rack signs.

Attachment 1: VacuumNormal.png
VacuumNormal.png
Attachment 2: vacValvesDisabled.jpg
vacValvesDisabled.jpg
  9016   Thu Aug 15 21:42:53 2013 CharlesUpdateISSISS - Schematic + PCB Layout

 After many, many moons of getting to know exactly how frustrating Altium can be, I have completed the PCB layout for my ISS board (final page of ISS_v3.pdf).

Before I get into detail about the PCB, there is one significant schematic change to note: the comparator circuit was changed (with significant help from Koji) so that the voltage reference for boost triggering is established in a more logical way. Instead of the somewhat convoluted topology I had before, now there are only two feedback resistors, R82 and R83. Because their resistances (500k and 50k respectively) are so much larger than the total resistance of the 1k potentiometer (used to establish a tunable threshold voltage), the current flowing through the feedback loop is negligible compared to the 5 mA current flowing through the potentiometer (the pot is rated for 2 W and with 5 mA -> 25 mW dissapation). This allows one to set the threshold voltage for my schmitt trigger, at pin 2 of both the pot and the comparator, entirely with the pot. This trigger also has hysteresis given by the relation deltaV ~ (R83/R82) * (Voh - Vol) where deltaV is the separation between threshold voltages, Voh is the high-level comparator ouput and Vol is the low-level comparator output. Koji simulated this using CircuitLab and I plan to verify the behavior by making a quick prototype circuit.

Now, on to the PCB. The board itself is of a 'standard' LIGO size (11" x 6") has 3 routing layers and 3 internal planes, one for +15 V, one for -15 V and one for GND. In the attached pdf, red is the top routing layer, blue is the bottom layer and brown is the middle routing layer (used for ±5 V exclusively). The grey circles are pads and vias (drilled through) and anything in black is silkscreen overlay. I placed each component and track by hand, attempting to minimize the signal path and following the general rules below,

  • Headers for power, ±5 V and ±15V, are at the back of the board
  • For sections of the board such as filter stages or buffers, resistors and capacitors were grouped around their respective op-amps.
  • As often as was possible, routing was confined to the top layer. Tracks on the bottom layer were placed mostly out of necessity (i.e. no possible connection on top routing layer).
  • The signal generally proceeds from left to right (directions with respect to the attached printout) in the same logical order as on the schematic sheets. Refer to the global sheet (page 1) of the attached "ISS_v3.pdf".
  • External ports such as the PD input, various monitoring ports and panel mounted switches/LEDs were all connected to the board via headers located along the front edge. These are also ordered following the schematic layout.
  • Occasionally, similar signal paths were grouped together although this was a rarity on my board

Sections of the board have been partitioned and labeled with silkscreen overlay to help in both signal pathway recognition as well as eventual troubleshooting.

On the board, I have also included holes so that it can be mounted inside of an enclosure. There is a DCC number printed as well as a 'barcode' (TrueType font: IDAutomationC39S), although they both contain filler asterisks as I haven't published this to the DCC and thus do not have a number.

Attachment 1: ISS_v3.pdf
ISS_v3.pdf ISS_v3.pdf ISS_v3.pdf ISS_v3.pdf ISS_v3.pdf ISS_v3.pdf
Attachment 2: ISS_v3-Power_Reg.pdf
ISS_v3-Power_Reg.pdf
  9015   Thu Aug 15 19:05:07 2013 manasaUpdateGreen LockingALS out of loop noise

Beat notes were recovered for both the arms.

I locked the arms to IR using PDH and measured the ALS out of loop noise at the phase tracker output.

The Y arm has the same 300Hz/rtHz rms. The X arm rms noise measures nearly the same as the Y arm in the 5-500Hz region (X arm has improved nearly 10 times after the last whitening filter stage change  old elog ).

The noise in the ALSX error signals could be related to the bad alignment and conditions at the X end.

Attachment 1: ALS_OutLoop.pdf
ALS_OutLoop.pdf
  9014   Thu Aug 15 12:30:17 2013 manasaUpdateGreen LockingLost beat notes

[Koji, Nic, Manasa]

Update from last night.

Koji and I realigned the green optics on the PSL to start working on the ALS.

We set on a beat note search. We couldn't find the beat note between any of the arm green transmission and the PSL green. All we could see was the beat between the X arm and the Y arm green leakage.

Since we had the beatnote between the 2 green transmission beams, we decided to scan the PSl temperature. We scanned the SLOW actuator adjust of PSL; but couldn't locate any beat note. The search will continue again today.

  9013   Thu Aug 15 09:34:12 2013 SteveUpdateGeneralWilcoxon cables rescued

Eric and Steve,

 We removed Wilcoxon Accelerometer PS and Amplifier unit under the BS optical tabel yesterday. The six cabels going to DAQ  were labeled and left in place. Gain setting were 100, except channel 3 was 10.

The ~ 40 m long 2 sets of 3 cables were very happy to get their kinks out. Especially the set going just south of ITMX optical  table.

We have to take better care of these cables! Your data will be useless this way.

Attachment 1: rescuedGraycables.jpg
rescuedGraycables.jpg
Attachment 2: wilconoxOut.jpg
wilconoxOut.jpg
Attachment 3: chanGains.jpg
chanGains.jpg
  9012   Thu Aug 15 01:51:50 2013 KojiSummaryGeneralRFM<->Dolphin bridge distributed to c1rfm and c1mcs

Since the RFM-Dolphin bridges for the ASX model was added to the c1rfm model, c1rfm kept timing-out from the single sample time of 60us.

The model had 19 dolphin accesses, 21 RFM accesses, and 9 shared memory (SHM) accesses.

At the beginning 2 RFM and 2 SHM accesses were moved to c1sus (i.e. they were mistakenly placed on c1rfm).
But this actually made the c1sus model timed out. So the model was reverted.

The current configuration is that the WFS related bridges were accommdated in the c1mcs model.
This made the timing of c1rfm ~40us. So it is safe now.
On the other hand, the c1mcs model has the time consumption of ~59us. This is marginal now.

We need to understand why any RFM access takes such huge delay.

  9011   Wed Aug 14 08:24:20 2013 SteveUpdateVACpumpdown at day 8

 

 

Attachment 1: pd76md8.png
pd76md8.png
  9010   Tue Aug 13 22:21:12 2013 KojiSummaryGeneralMinicircuit Filter TFs (AG4395A test)

As a part of the network analyzer test in the previous entry, the transfer functions of Mini-Circuits filters we have at the 40m were measured.

<<List of the filters>>

- LPF (SMA): SLP1.9, SLP5, SLP21.4, SLP30, SLP50, SLP100, SLP150, SLP750
- LPF (BNC): BLP1.9, BLP2_5, BLP5, BLP30
- BPF (SMA): SBP10.7, SBP21.4, SBP70
- HPF (SMA): SHP25, SHP100, SHP150, SHP200, SHP500

 

Attachment 1: Minicircuit_LPF.pdf
Minicircuit_LPF.pdf
Attachment 2: Minicircuit_BPF.pdf
Minicircuit_BPF.pdf
Attachment 3: Minicircuit_HPF.pdf
Minicircuit_HPF.pdf
Attachment 4: 130813.zip
  9009   Tue Aug 13 21:49:32 2013 KojiSummaryGeneralTesting new AG4395A network analyzer

New AG4395, sn MY41101114  for West Bridge Labs was delivered. For the test purpose it is at the 40m now.

I made a series of tests in order to find anything broken.

Network analyzer test

- RF out / Rch test

RF out directly connected to R input channel.
The received power at the R-ch was measured while the output was swept from 10Hz to 500MHz.

The RF power was changed from -50dBm to +15dBm with +10dBm increment (but the last one).

The attenuator setting was changed from 50dB to 0dB.

=> The configured output power was properly detected by the R channel.

=> RF output is producing the signal properly. R-ch is detecting the produced signal properly.

- Ach/Bch test

Same test as above for Ach and Bch 

=> Same result as above

=> A-ch and B-ch are detecting the produced signal properly.

- Transfer function test

Connect a power splitter to the RF out. Detect the split signals by R-ch and A-ch

=> Measurement is at around 0dB +/- 1dB up to 500MHz.

Same measurement for B-ch

=> Same result

=> A/R and B/R indicates proper transfer function measurements.

- Calibration

RF out was split in to two. One was connected to R-ch. The other was connected to A-ch.
The thru response calibration was run.

=> The thru calibration was performed properly. 

- Practical tranfer function measurements.

In the above calibration setup, various RF filters were inserted in the Ach path.

The measured data was extracted via GPIB connection.

=> Practical transfer function measurements were performed.

=> GPIB connectivity was confirmed

 

External reference test

- External 10MHz reference from an SRS frequency counter was connected to Ext Ref In

=> Ext Ref indicator on the screen appeard

=> The internal oscillator seemed to be locked to the external reference in

 

 

Spectrum analyzer test

- Measured the signals from DS345 by R/A/B ch

Sinusoidal signal (1V) swept from 10MHz to 30Mhz

=> Corresponding moving peak was detected in each case

- Noise level measurement

R/A/B channels were terminated. The attenuation at each port was set to 0dB.

Frequency span was changed between 500MHz, 10MHz, 100kHz, 1kHz.

=> Noise level of ~10nV/rtHz between 0.1-500MHz was confirmed. All R/A/B channels have the same performance.

Attachment 1: AG4395A_noise.pdf
AG4395A_noise.pdf
  9008   Tue Aug 13 21:09:03 2013 manasaUpdateGreen LockingArms ready for ALS

I aligned both the X and Y end green to the arms.

The transmitted green were aligned at the PSL table green optics to the beat PDs.
Beat notes were retrieved.
 
To do:
1. Check Y arm ALS with previous performance.
2. Troubleshoot X arm ALS.
3. Edit the automation scripts for ALS.
4. Modify ALS model to talk to LSC instead of suspension models.
  9007   Tue Aug 13 17:20:54 2013 KojiUpdateCDS[Fixed] c1iscex needs help

c1x01 timing issue was solved. Now all of the models on c1iscex are nicely running.

Symptons

- c1x01 was synchronized to 1PPS in stead of TDS

- C1:DAQ-DC0_C1X01_STATUS (Upper right indicator) was red. The bits were 0x4000 or 0x2bad.
  C1:DAQ-DC0_C1X01_CRC_SUM kept increasing

 - c1scx, c1spx, c1asx could not get started.

Solution

- login to c1iscex "ssh c1iscex"

- Run "sudo shutdown -h now"

- Walk down to the x end rack

- Make sure the supply voltages for the electronics are correct (See Steve's entry)

- Make sure the machine is already shutdown.

- Unplug two AC power supply of the machine.

- Turn off the front panel switch of the IO chassis

- Wait for 10sec

- Turn on the IO chassis

- Plug the AC power supply cables to the machine

- Push the power switch of the realtime machine

  9006   Tue Aug 13 13:30:41 2013 Alex ColeConfigurationElectronicsCable Routing

 I routed cables (RG405 SMA-SMA) from several demodulator boards in rack 1Y2 to the RF Switch in rack 1Y1 using the overhead track. Our switch chassis contains two 8x1 switches. The COM of the "right" switch goes to channel 7 of the "left" switch to effectively form a 16x1 switch. The following is a table of correspondences between PD and RF Switch input.

 

PD Left/Right Switch Channel Number
REFL11

R

0
POX11 L 0
AS55 R 1
REFL55 R 7
POP22 R 6
REFL165 R 5
REFL33 L 7

 

ThePOP110 demod board has not yet had a cable routed from it to the switch because I ran out of RG405.

We should also consider how important it is to include MCREFL in our setup. Doing so would require fabrication of a ~70 ft RG405 cable. 

Attachment 1: photo_(6).JPG
photo_(6).JPG
  9005   Tue Aug 13 11:54:40 2013 Alex ColeHowToElectronicsRF PD Fiber-Coupled Laser Operation

This post pertains to the fiber-coupled diode laser mounted in rack 1Y1.

To turn the laser on, first turn the power supply's key (red) to the clockwise. Then make sure that the laser is in "current" mode by checking that the LED next to "I" in the "Laser Mode" box in lit up. If the light is not on, press the button to the right of the "I" light until it is. Now press the output button (green). This is like removing the safety for the laser. Then turn the dial (blue) until you have your desired current. Presently, the current limit is set to around 92 mA.

To turn the laser off, dial the current back down to 0mA and turn the key (red) counterclockwise.

Attachment 1: photo_(4).pdf
photo_(4).pdf
  9004   Tue Aug 13 11:40:19 2013 Alex ColeSummaryElectronicsRFPD Demod Filter Frequency Response Measurement

 For the RF PD Frequency Response Measurement project, we get each PD signal from the "PD RF Mon" output of each demodulator board corresponding to our PD under test. Therefore we can't neglect the frequency response of various filters inside the demodulator board. I used our Agilent 4395 Network Analyzer to gather frequency response data for each demodulator board being considered for the RFPD frequency response project (AS55, REFL11, REFL33, REFL55, REFL165, POX11, POP22, POP110).

The NA swept over a frequency range of 1-500 MHz. Data was collected using NWAG4395A (from the netgpibdata directory). It should be noted that the command line options -a 16 -x 15 (averaging=16 and excitation amplitude=15 dBm[the max]), in addition to the usual command line options described in the help file, were used to minimize noise. 

The data is located in /users/alex.cole. The file names are in the format [PDNAME]DemodFilt_1000000.dat (e.g. REFL11DemodFilt_1000000.dat). Results for POP110 are shown below.

Attachment 1: photo_(3).JPG
photo_(3).JPG
Attachment 2: test.jpg
test.jpg
  9003   Tue Aug 13 11:04:44 2013 SteveUpdatePEMfluorecent lights

Our fluorecent lights became obsolete.  We'll have change fixtures over to some more energy efficient one. Do you have any recommendation regarding to less noise performer unit?

We may go this direction of LED fluorecent lamps ?

  9002   Tue Aug 13 07:40:53 2013 SteveUpdateCDSc1iscex needs help

 

 Sorrensen ps ouput of +15V at rack 1X9 was current limited to 10.3V @ 2A

Increased threshold to 2.1A  and the voltage is up to 14.7V

Attachment 1: c1iscexSick.png
c1iscexSick.png
  9001   Mon Aug 12 23:13:14 2013 JenneUpdateASCPRCL ASS software in place

Quote:

Why POPDC???

 I guess I was thinking that POPDC was a proxy for any type of PRCL lock.  Even if we're sideband locked, there is still some signal in POPDC (although it is very small relative to a carrier lock - ~40cts vs. 1,000cts).  However, as soon as this question was asked of me, I realized that one of the 2f demodulated signals made more sense. 

Since I want the ability to choose between POP110 and POP22, I have put a little 1x3 input matrix before the PRCL lockins in the ASS model.  Since POPDC was already there, I included it as an option in the matrix (in case we ever want to do some PRCL ASS after we have some carrier resonating as well). 

  9000   Mon Aug 12 21:27:03 2013 manasaUpdateCDSc1iscex needs help

I started to modify the c1asx model to reduce the RFM model from hitting its max time.
Instead of bringing in ASS, I have modified ASX to do everything and only the clock signals to ITMX pitch and yaw are now going through RFM. RFM is still hitting 62usec and I suppose that is because of the problems with c1iscex.

c1iscex not happy

Cause and symptoms

While restarting the models, c1iscex crashed a couple of times because of some errors and had to be powercycled. The models were modified and they seem to start ok.
But it looks like there is something wrong with c1iscex since the models were started. The GPS time is off and C1:DAQ-DC0_C1X01_CRC_SUM keeps building up even for c1x01 which was left untouched.

Trial treatments

1. Since c1x01 ans c1spx were not touched,c1scx and c1asx were killed and we tried to start the other models. This did not help.
2. Koji did a manual daqd restart which did not help either.

We are leaving c1iscex as is for the time being and calling Jamie for help.

P.S. While making the models, I had created IPCx_PCIE blocks in c1iscex which do not exist. I changed them to RFM and SHMEM blocks. This did not allow me to compile the model and was only spitting errors of IPCx mismatch. After some struggle and elog search I figured out from an old elog that eventhough the IPCx blocks are changed in the model, the old junk exists in the ipc file in chans directory. I deleted all junk channels related to the ASX model. The model compiled right away.

  8999   Mon Aug 12 17:30:03 2013 KojiUpdateASCPRCL ASS software in place

Why POPDC???

  8998   Mon Aug 12 15:39:40 2013 JenneUpdateASCPRCL ASS screens in place

I have added the PRCL ASS to the main ASS screen, and created the servo and lockin screens.  The filters loaded are the same as those used for the arms (bandpasses and lowpasses for the lockins, and an integrator for the servo).

I'm going to try to lock, and get the ASS to work.

  8997   Mon Aug 12 14:05:34 2013 JenneUpdateASCPRCL ASS software in place

Quote:

- We are ready to implement ASS for PRM

 I have added an IPC sender from the LSC model, to send POPDC to ASS.  I have copied over the structure of the arms' ASS, to do the same for PRCL.  I have set it up to dither the PRM, and feed back to the PRM.  I did not include an LSC set, since I'm assuming that we'll set the input pointing with the arms, and just want to move the PRM to maximize POPDC.

Models have been compiled, installed, and restarted, and the daqd was restarted.

  8996   Mon Aug 12 13:30:33 2013 JamieUpdateCDSX-End Green ASS - Roundup

Quote:

I'm not really sure why the ASS was involved in this.  I feel like it might have been simpler to just do everything in the ASX model, to keep things cleaner.  Also, the IPC blocks for this stuff (in both ASS and ASX) are not on the top level of the model.  I had thought that this was expressly forbidden (although I'm not sure why).  I'm emailing Jamie, to see if he remembers what, if anything, is breakable if the IPC blocks are down a level.

I'm not sure if it's forbidden by the RCG, but you should definitely NOT do it.  All IO, whether it be between ADC/DACs or IPCs, should always be at the model top level.  That's what keeps things portable, and makes it easier to keep track of where are signals are going/coming from.

  8995   Mon Aug 12 12:57:59 2013 JenneUpdateCDSX-End Green ASS - Roundup

Quote:
  1. The SIMULINK model has been modified to accommodate an option to dither the cavity mirrors and not the PZT mirrors. Details are as follows:
    • I have sent the LO signals (CLK,SIN and COS) from the ASS model to the ASX model via the RFM model. Appropriate changes were made to all these three models, and recompiling and restarting the models was done without issue. The SIN and COS signals are used to demodulate green transmission at the dither frequencies. ***The CLK signal is not required to be sent between models as it is not being used by ASX (I turn the dither ON using the channels already set up for ASS). I realised this a little late, and at present the ASS and RFM models are compiled such that the CLK signal is also sent from ASS to RFM. This can be removed, thus freeing up 4 unnecessary inter-process communication channels. Also, I am not too sure if this is relevant, but the maximum computation time of both the RFM and ASX models seem to have gone up after I added these inter-process communication links.***

    •  

 Getting rid of the LO transmission will certainly help / be good.  After adding these channels, the RFM model is regularly hitting 62usec (out of a max acceptable of 60).

I'm not really sure why the ASS was involved in this.  I feel like it might have been simpler to just do everything in the ASX model, to keep things cleaner.  Also, the IPC blocks for this stuff (in both ASS and ASX) are not on the top level of the model.  I had thought that this was expressly forbidden (although I'm not sure why).  I'm emailing Jamie, to see if he remembers what, if anything, is breakable if the IPC blocks are down a level.

ELOG V3.1.3-